# ADBM: Adversarial Diffusion Bridge Model for Reliable Adversarial Purification

### v1

@@ 5.2 Robustness against White-Box Adaptive Attacks
We first evaluate ADBM against the reliable while-box adaptive attacks to show the worst-case adversarial robustness where the attacker has complete knowledge. Note that we expect a defense method to be robust not only on seen threats but also on unseen attack threats. Thus, unless otherwise specified, we evaluated the models on three attack threats: $l_{\infty}$, $l_{1}$, and $l_{2}$, with the bounds $\epsilon_{\infty}=8/255$, $\epsilon_{1}=12$, $\epsilon_{2}=1$, respectively. Here $l_{\infty}$ attack is considered the seen threat as ADBM was trained with $l_{\infty}$ adversarial noise, while $l_{1}$ and $l_{2}$ attacks can be regarded as unseen threats.
We compared ADBM with SOTA AT and AP methods. All AT models were trained also with $l_{\infty}$ adversarial examples only, ensuring $l_{1}$ and $l_{2}$ threats were unseen for these models. We used AutoAttack [32] to implement the adversarial threats for AT methods, while we used the attack practice described in Sec. 3 (PGD with 200 iteration steps and 20 EOT samples) to implement the adversarial threats for AP methods. Additional configurations of these attacks can be found in Sec. C.2. Note that this evaluation comparison is justified since Lee and Kim [19] observed that AutoAttack with EOT yielded inferior attack performance compared to PGD with EOT for stochastic pre-processing defenses. Our experiments in Sec. D.1 further confirmed that PGD with EOT is the best practice than other attacks such as AutoAttack, C&W, and DeepFool with EOT.
The accuracies under the reliable white-box adaptive attacks on CIFAR-10 are shown in Tab. 3. In this evaluation, we compared ADBM with several competitive methods. Both Gowal et al. [24] and Wang et al. [25] performed AT with about 50M generated images by diffusion models. Additionally, we compared ADBM with several AP methods using generative models, especially diffusion models. By inspecting these results, we can see that despite training with millions of examples, AT methods still exhibited limited robustness against unseen $l_{1}$ and $l_{2}$ attacks. In contrast, ADBM demonstrates a strong defense against $l_{1}$ and $l_{2}$ attacks. Moreover, attempts to enhance the performance of DiffPure by introducing input guidance [26] or applying Langevin dynamics [27] occasionally resulted in detrimental effects when assessed through reliable adaptive attack evaluations. Notably, ADBM outperformed DiffPure by achieving an average robustness gain of 4.4% on CIFAR-10, while the clean accuracies kept comparable. Similar outcomes were observed for SVHN, Tiny-ImageNet, and ImageNet-100, detailed in Sec. D.2, reinforcing similar findings from the CIFAR-10 analyses.
@@
WRN-28-10 [43] AT $94.46$ $68.90$ $58.96$ $74.22$ $87.60$ $76.83$
WRN-28-10 [44] $93.00$ $62.30$ $60.49$ $71.78$ $87.60$ $75.53$
WRN-28-10 [25] $95.56$ $75.16$ $67.23$ $80.39$ $88.57$ $81.38$
UNet+WRN-28-10 DiffPure [16] AP $93.93$ $92.97$ $92.15$ $92.19$ $91.49$ $92.20$
@@ 5.3 Robustness against Black-Box Attacks
We then considered the more realistic black-box attacks, where the attacker has no knowledge about the defense mechanism, i.e., the purification model, and cannot access the gradients of models. Instead, the attacker can only query the model’s output with query-based attacks or use substitute models with transfer-based attacks. We conducted three query-based attacks: Square [45], RayS [46], and SPSA [47] attacks. Square and RayS are efficient search-based attack methods, while SPSA approximates gradients by sampling randomly in a black-box manner. For Square and RayS, we utilized 5,000 search steps. In the SPSA attack, we set $\sigma$ to 0.001, the number of random samples in each iteration step to 128, and the iteration step to 40. The transfer-based attack was implemented by generating adversarial examples via the vanilla models.
The results under black-box attacks on SVHN and CIFAR-10 are shown in Tab. 4 and Sec. D.3, respectively. We can see that these black-box attacks achieved excellent attack performance on the vanilla model. For the SOTA AT method [25] on SVHN, these attacks lowered the average accuracy to 81.38%. But surprisingly, all these black-box attacks can hardly lower the accuracies of ADBM. Thus, we can conclude that besides the promising results on white-box attacks, under the realistic black-box attacks, ADBM has advantages over AT models even on the seen threat (i.e., $l_{\infty}$ threat).
@@
5.4 Ablation Study
Reverse steps. We first investigated the influence of reverse steps on the adversarial robustness of ADBM. The number of reverse steps is proportional to the inference cost. We used five reverse steps in the main experiments. Here we evaluated the robustness of ADBM with fewer steps to investigate whether the number of steps can be further reduced. The results in Tab. 6 show ADBM was more robust than DiffPure regardless of reverse steps. Notably, even with just one reverse step, ADBM maintained its good robustness. We discuss the detailed inference cost related to this further in Sec. 6.
Effectiveness on new classifiers. To assess the effectiveness of ADBM on new classifiers, we conducted a study to investigate its transferability. Specifically, we utilized the fine-tuned ADBM checkpoint, trained with adversarial noise from a WRN-70-16 classifier, as the pre-processor for a WRN-28-10 model and a vision transformer model [48] directly, denoted as ADBM(Transfer). The results shown in Sec. D.4 demonstrate that ADBM(Transfer) achieved robust accuracies comparable to ADBM directly trained with corresponding classifiers. This finding highlights the practicality of ADBM, as the fine-tuned ADBM model on a specific classifier can potentially be directly applied to a new classifier without requiring retraining. We guess this may be attributed to the transferability of adversarial noise [49].
@@ 6 Conclusion and Discussion
In this work, we introduce ADBM, a cutting-edge method for diffusion-based adversarial purification. Theoretical analysis supports the superiority of ADBM in enhancing adversarial robustness. With extensive experiments, we demonstrated the effectiveness of ADBM across various scenarios using reliable adaptive attacks. Notably, ADBM demonstrates significant improvements over prevalent AT methods against unseen threats or black-box attacks. In the era of foundation models [50], training foundation models with AT becomes increasingly challenging due to high computational costs. ADBM, on the other hand, provides a promising alternative as a plug-and-play component that can enhance the robustness of existing foundation models without the burdensome of retraining. The social impact of this work is discussed in Appendix E.
Limitation. While ADBM requires quite few reverse steps, it does introduce additional computational demands due to its reliance on diffusion models. The current UNet architectures, tailored primarily for generative purposes, are quite large (see Sec. D.5). Exploring the potential of downsizing these architectures for AP remains an open area for future research.
Appendix A Evaluations on DiffPure
A.1 Previous Evaluation Methods on DiffPure
@@
As the original implementation of DiffPure needs dozens of prediction steps in the reverse process, it is challenging to compute the full gradient of the whole purification process due to memory constraints. To circumvent computing the full gradient, Nie et al. [16] originally employed the adjoint method (along with Expectation over Transformation (EOT) [5]) to compute an approximate gradient. But recent works have identified it to be a flawed adaptive attack [19, 20]. As improved adaptive mechanisms, Lee and Kim [19] proposed a surrogate attack for DiffPure (to approximate the gradient of the iterative reverse procedure better) and Chen et al. [20] utilized the gradient-checkpointing technique [51] to trade time for memory space and compute the full gradient directly. Despite previous efforts, we find that the adaptive evaluations for DiffPure remained insufficient. Specifically, the surrogate attack failed to compute the full gradient, and the gradient-checkpointing attack in Chen et al. [20] employed insufficient iteration steps or EOT samples, an issue underscored by Gao et al. [52] which highlights the importance of adequate steps and samples for evaluating stochastic pre-processing defenses like DiffPure.
A.2 The Cost of Our Evaluation
@@
In our evaluation, we used the full gradient of the whole reverse process and set the PGD iteration steps to 200, with 20 EOT samples for each iteration. The cost of the evaluation is quite high, especially in the context of the high reverse steps of the original DiffPure. To give a concrete example, for a single input image, the noise prediction model (i.e., ${\boldsymbol{\epsilon}}_{\theta}$) in the original DiffPure implementation with 100 reverse steps needs to be queried a total of $400,000$ times.
@@
However, we think that such high efforts are worthwhile as an unreliable evaluation could create a false sense of security on defenses. Historical evidence has shown that many defenses initially considered robust, were subsequently breached by more sophisticated and dedicated attacks [5, 6]. Our aim is to prevent a similar outcome for diffusion-based purification, advocating for the employment of a meticulous and reliable attack evaluation methodology, regardless of the expense involved.
@@
On the other hand, with our reliable evaluation, we have investigated the influence of forward steps and reverse steps on the robustness of diffusion-based purification. Our following results in Tab. 2 indicate that reverse steps will not significantly influence the robustness of DiffPure under the reliable evaluation. This suggests that diffusion-based purification techniques might not benefit from increasing the number of reverse steps to complicate the attack process, as such strategies could finally be neutralized by high-cost attacks similar to the one we have implemented. And when the reverse steps are reduced (e.g., five steps in our main experiments), the attack cost of our evaluation method is comparable to the most widely used AutoAttack benchmark [32].
A.3 Additional Investigation on DiffPure with Our Evaluation
@@
Appendix B Derivation of Equations and Proofs of Theorems
B.1 The Derivation of Eq. 7
@@ Appendix B Derivation of Equations and Proofs of Theorems
Since ${\mathbf{x}}_{t}^{d}={\mathbf{x}}_{t}^{a}-k_{t}{\boldsymbol{\epsilon}}_{a}$ and ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$, there is
${\mathbf{x}}_{t}^{d}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}.$ (13)
@@
$q({\mathbf{x}}_{t-1}^{d}|{\mathbf{x}}_{t}^{d},{\mathbf{x}}_{0})$ $=\frac{q({\mathbf{x}}_{t}^{d}|{\mathbf{x}}_{t-1}^{d},{\mathbf{x}}_{0})\cdot q({\mathbf{x}}_{t-1}^{d}|{\mathbf{x}}_{0})}{q({\mathbf{x}}_{t}^{d}|{\mathbf{x}}_{0})}\propto\exp(-\frac{1}{2}(\frac{({\mathbf{x}}_{t}^{d}-\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1}^{d}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\boldsymbol{\epsilon}}_{a})^{2}}{1-\alpha_{t}}$ (14)
$+\frac{({\mathbf{x}}_{t-1}^{d}-\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\boldsymbol{\epsilon}}_{a})^{2}}{1-\bar{\alpha}_{t-1}}-\frac{({\mathbf{x}}_{t}^{d}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a})^{2}}{1-\bar{\alpha}_{t}}))$
$=\exp\left(-\frac{1}{2}(A({\mathbf{x}}_{t-1}^{d})^{2}+B{\mathbf{x}}_{t-1}^{d}+C({\mathbf{x}}_{0},{\boldsymbol{\epsilon}}_{a},{\mathbf{x}}_{t}^{d}))\right),$
@@
$B=-2\sqrt{\alpha_{t}}\cdot\frac{{\mathbf{x}}_{t}^{d}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\boldsymbol{\epsilon}}_{a}}{1-\alpha_{t}}-2\frac{\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}+(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\boldsymbol{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}.$
In inference, we expect that ${\boldsymbol{\epsilon}}_{a}$ in Eq. 14 can be eliminated since only ${\mathbf{x}}_{0}^{a}$ is given by attackers and ${\boldsymbol{\epsilon}}_{a}$ cannot be decoupled from ${\mathbf{x}}_{0}^{a}$ directly. Based on the property of Gaussian distribution, eliminating all terms related to ${\boldsymbol{\epsilon}}_{a}$ in $B$ and $C$ in Eq. 5 can be achieved by eliminating all terms related to ${\boldsymbol{\epsilon}}_{a}$ in $B$. This yields:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\boldsymbol{\epsilon}}_{a}}{1-\alpha_{t}}-\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\boldsymbol{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}=0.$ (16)

### v3

@@ 5.2 Robustness against White-Box Adaptive Attacks
We first evaluate ADBM against the reliable while-box adaptive attacks (Athalye et al., 2018) to show the worst-case adversarial robustness where the attacker has complete knowledge. Note that we expect a defense method to be robust not only on seen threats but also on unseen attack threats. Thus, unless otherwise specified, we evaluated the models on three attack threats: $l_{\infty}$, $l_{1}$, and $l_{2}$, with the bounds $\epsilon_{\infty}=8/255$, $\epsilon_{1}=12$, $\epsilon_{2}=1$, respectively. Here $l_{\infty}$ attack is considered the seen threat as ADBM was trained with $l_{\infty}$ adversarial noise, while $l_{1}$ and $l_{2}$ attacks can be regarded as unseen threats.
We compared ADBM with SOTA AT and AP methods. All AT models were trained also with $l_{\infty}$ adversarial examples only, ensuring $l_{1}$ and $l_{2}$ threats were unseen for these models. We used AutoAttack (Croce & Hein, 2020), which is recognized as the strongest attacks for AT methods, to implement the adversarial threats for AT methods, while we used the attack practice described in Sec. 3 (PGD with 200 iteration steps and 20 EOT samples) to implement the adversarial threats for AP methods. Additional configurations of these attacks can be found in Sec. D.2. Note that this evaluation comparison is justified since Lee & Kim (2023) observed that AutoAttack with EOT yielded inferior attack performance compared to PGD with EOT for stochastic pre-processing defenses. Our experiments in Sec. E.1 further confirmed that PGD with EOT is the best practice than other attacks such as AutoAttack, C&W, and DeepFool with EOT.
The accuracies under the reliable white-box adaptive attacks on CIFAR-10 are shown in Tab. 3. In this evaluation, we compared ADBM with several competitive methods. Both Gowal et al. (2021) and Wang et al. (2023) performed AT with about 50M generated images by diffusion models. Additionally, we compared ADBM with several recent AP methods using generative models, especially diffusion models. [^3] RDC (Chen et al., 2023a) was not compared as it required thousands of times the inference cost of DiffPure so that we cannot afford to conduct experiments in our setting. Instead, we discuss it in Sec. A.2. By inspecting these results, we can see that despite training with millions of examples, AT methods still exhibited limited robustness against unseen $l_{1}$ and $l_{2}$ attacks. In contrast, ADBM demonstrates a strong defense against $l_{1}$ and $l_{2}$ attacks. Moreover, attempts to enhance the performance of DiffPure by introducing input guidance (DiffPure+Guide, Wang et al. (2022)) or applying Langevin dynamics (DiffPure+Langevin, Zhang et al. (2023)) occasionally resulted in detrimental effects when assessed through reliable adaptive attack evaluations. Notably, ADBM outperformed DiffPure by achieving an average robustness gain of 4.4% on CIFAR-10, while the clean accuracies kept comparable. Similar outcomes were observed for SVHN, Tiny-ImageNet, and ImageNet-100, detailed in Sec. E.2, reinforcing similar findings from the CIFAR-10 analyses.
@@
WRN-28-10 (Rade et al., 2022) AT $94.46$ $68.90$ $58.96$ $74.22$ $87.60$ $76.83$
WRN-28-10 (Yang et al., 2023) $93.00$ $62.30$ $60.49$ $71.78$ $87.60$ $75.53$
WRN-28-10 (Wang et al., 2023) $95.56$ $75.16$ $67.23$ $80.39$ $88.57$ $81.38$
UNet+WRN-28-10 DiffPure (Nie et al., 2022) AP $93.93$ $92.97$ $92.15$ $92.19$ $91.49$ $92.20$
@@ 5.3 Robustness against Black-Box Attacks
We then considered the more realistic black-box attacks, where the attacker has no knowledge about the defense mechanism, i.e., the purification model, and cannot access the gradients of models. Instead, the attacker can only query the model’s output with query-based attacks or use substitute models with transfer-based attacks.
On the seen threat (i.e., $l_{\infty}$ threat), we conducted three query-based attacks: RayS (Chen & Gu, 2020), Square (Andriushchenko et al., 2020), and SPSA (Uesato et al., 2018) attacks. Square and RayS are efficient search-based attack methods, while SPSA approximates gradients in a black-box manner. For Square and RayS, we utilized 5,000 search steps. For SPSA, we set $\sigma$ to 0.001, the number of random samples in each step to 128, and the iteration step to 40. In addition, We implemented a transfer-based attack by generating adversarial examples via the vanilla models. The results under black-box attacks on SVHN and CIFAR-10 are shown in Tab. 4 and Sec. E.3, respectively. We can see that these black-box attacks achieved excellent attack performance on the vanilla model. For the SOTA AT method on SVHN, these attacks lowered the average accuracy to 81.38%. But surprisingly, all these black-box attacks can hardly lower the accuracies of ADBM. Thus, we can conclude that besides the promising results on white-box attacks, under the realistic black-box attacks, ADBM has advantages over AT models even on the seen threat.
We further conducted experiments on more unseen threats beyond norm-bounded attacks, including patch-like attacks (Gao et al., 2020; Yuan et al., 2022) and recent diffusion-based attacks (Chen et al., 2023b; Xue et al., 2023), to evaluate the generalization ability of ADBM. The results shown in Sec. E.4 indicate that ADBM demonstrates better generalization ability than DiffPure and AT methods against these black-box unseen threats.
5.4 Ablation Study
Reverse sampling steps. We first investigated the influence of reverse sampling steps on the adversarial robustness of ADBM. The number of reverse steps is proportional to the inference cost. We used five reverse steps in the main experiments. Here we evaluated the robustness of ADBM with fewer steps to investigate whether the number of steps can be further reduced. The results in Tab. 6 show ADBM was more robust than DiffPure regardless of reverse steps. Notably, even with just one reverse step, ADBM maintained its good robustness. We discuss the detailed inference cost related to this further in Sec. 6.
@@ 5.4 Ablation Study
Effectiveness on new classifiers. To assess the effectiveness of ADBM on new classifiers, we conducted a study to investigate its transferability. Specifically, we utilized the fine-tuned ADBM checkpoint, trained with adversarial noise from a WRN-70-16 classifier, as the pre-processor for a WRN-28-10 model and a vision transformer model (Dosovitskiy et al., 2021) directly, denoted as ADBM (Transfer). The results shown in Sec. E.5 demonstrate that ADBM (Transfer) achieved robust accuracies comparable to ADBM directly trained with corresponding classifiers. This finding highlights the practicality of ADBM, as the fine-tuned ADBM model on a specific classifier can potentially be directly applied to a new classifier without requiring retraining. We guess this may be attributed to the transferability of adversarial noise (Dong et al., 2018).
@@ 6 Conclusion and Discussion
In this work, we introduce ADBM, a cutting-edge method for diffusion-based adversarial purification. Theoretical analysis supports the superiority of ADBM in enhancing adversarial robustness. With extensive experiments, we demonstrated the effectiveness of ADBM across various scenarios using reliable adaptive attacks. Notably, ADBM demonstrates significant improvements over prevalent AT methods against unseen threats or black-box attacks. In the era of foundation models (Bommasani et al., 2021), training foundation models with AT becomes increasingly challenging due to high computational costs. ADBM, on the other hand, provides a promising alternative as a plug-and-play component that can enhance the robustness of existing foundation models without the burdensome of retraining. In addition, our discussion on the acceleration of diffusion-based purification methods may unlock the potential of diffusion-based purification methods in various real-time applications (see Appendix F). The social impact of this work is further discussed in Appendix G.
Limitation. 1) While ADBM requires quite few reverse steps, it does introduce additional computational parameters due to its reliance on diffusion models compared with AT methods. The current UNet architectures, tailored primarily for generative purposes, are a bit large (see Appendix F). Exploring the potential of downsizing these architectures for AP remains an open area. 2) While ADBM requires only about 1/10 training steps of original diffusion models, it does introduce additional fine-tuning cost compared with using the pre-trained diffusion models directly (DiffPure). But we note that whether the slight extra training cost is worthwhile depends on the specific problems being addressed. In many security-critical cases, where the training cost is not the primary concern, ADBM offers an alternative approach to further enhance robustness with affordable training costs.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. U2341228).
Appendix A Discussions on Diffusion Bridges and Diffusion Classifiers
A.1 Diffusion Bridge Models
Bortoli et al. (2021) is fundamentally unsuitable for the AP task because it bridges the data distribution with a simple prior distribution, while Zhou et al. (2024); Bortoli et al. (2021); Liu et al. (2023b) tend to a training collapse for the AP task. Here we provide the theoretical explanation: Adversarial examples typically reside close to clean examples, particularly in the commonly analyzed $l_{\infty}$-bounded threat settings. This proximity poses a significant challenge to applying these bridge models (Zhou et al., 2024; Bortoli et al., 2021; Liu et al., 2023b) for AP, as these bridge models’ merely predicting the input adversarial image could result in extremely small training loss, finally incurring a degenerated solution and losing the capability to purify the adversarial perturbations.
Different from these bridge models, ADBM first maps the adversarial data distribution to a diffused (Gaussian) adversarial data distribution and then creates a bridge from the diffused adversarial data distribution to the clean data distribution. This is distinct from existing diffusion bridges, effectively solving the problems that previous diffusion bridges face in purification. Our theoretical analysis in Sec. 4.4 consolidates that ADBM guarantees superior performances for AP.
A.2 Robust Diffusion Classifiers
Chen et al. (2023a) show that a single diffusion model can be transformed into an adversarially robust diffusion classifier (RDC) using Bayes’ rule. Although RDC does not require an additional classifier, ADBM has two unique and significant advantages over RDC:
1) ADBM is an adversarial purification method, which can function as a preprocessing module in a plug-and-play manner without retraining the classifiers. This is a quite flexible and useful way, especially in the era of foundation models. For example, we may use ADBM as a preprocessing step to purify the visual inputs of a large vision-language model (VLM), without retraining the VLM with adversarial training (huge computational costs are saved). However, RDC does not follow this paradigm, as it is a robust classifier instead of a robust pre-processor. RDC is limited to generalize to tasks beyond robust classification, hindering its practical usefulness. Overall, ADBM offers a distinct advantage over RDC by decoupling preprocessing and the specific classification task, allowing each to be optimized independently.
2) While diffusion-based purification requires slightly more parameters (an additional classifier), ADBM requires significantly lower inference costs than RDC. Our following experiments indicate that, even with one reverse step, ADBM can maintain substantial robustness under reliable evaluation. In contrast, RDC requires N + T reverse steps, where T is the diffusion steps (e.g., 1000 in DDPM) and N is the number of likelihood maximization steps. The inference cost is proportional to the reverse steps, and thus the inference cost of RDC is several hundred times that of ADBM.
Appendix B Evaluations on DiffPure
B.1 Previous Evaluation Methods on DiffPure
@@
As the original implementation of DiffPure needs dozens of prediction steps in the reverse process, it is challenging to compute the full gradient of the whole purification process due to memory constraints. To circumvent computing the full gradient, Nie et al. (2022) originally employed the adjoint method (along with Expectation over Transformation (EOT) (Athalye et al., 2018)) to compute an approximate gradient. But recent works have identified it to be a flawed adaptive attack (Lee & Kim, 2023; Chen et al., 2023a). As improved adaptive mechanisms, Lee & Kim (2023) proposed a surrogate attack for DiffPure (to approximate the gradient of the iterative reverse procedure better) and Chen et al. (2023a) utilized the gradient-checkpointing technique (Chen et al., 2016) to trade time for memory space and compute the full gradient directly. Despite previous efforts, we find that the adaptive evaluations for DiffPure remained insufficient. Specifically, the surrogate attack failed to compute the full gradient, and the gradient-checkpointing attack in Chen et al. (2023a) employed insufficient iteration steps or EOT samples, an issue underscored by Gao et al. (2022) which highlights the importance of adequate steps and samples for evaluating stochastic pre-processing defenses like DiffPure.
B.2 The Cost of Our Evaluation
@@
In our evaluation, we used the full gradient of the whole reverse process and set the PGD iteration steps to 200, with 20 EOT samples for each iteration. The cost of the evaluation is quite high, especially in the context of the high reverse steps of the original DiffPure. To give a concrete example, for a single input image, the noise prediction model (i.e., ${\bm{\epsilon}}_{\theta}$) in the original DiffPure implementation with 100 reverse steps needs to be queried a total of $400,000$ times.
@@
However, we think that such high efforts are worthwhile as an unreliable evaluation could create a false sense of security on defenses. Historical evidence has shown that many defenses initially considered robust, were subsequently breached by more sophisticated and dedicated attacks (Athalye et al., 2018; Tramèr et al., 2020). Our aim is to prevent a similar outcome for diffusion-based purification, advocating for the employment of a meticulous and reliable attack evaluation methodology, regardless of the expense involved.
@@
On the other hand, with our reliable evaluation, we have investigated the influence of forward steps and reverse steps on the robustness of diffusion-based purification. Our following results in Tab. 2 indicate that reverse steps will not significantly influence the robustness of DiffPure under the reliable evaluation. This suggests that diffusion-based purification techniques might not benefit from increasing the number of reverse steps to complicate the attack process, as such strategies could finally be neutralized by high-cost attacks similar to the one we have implemented. And when the reverse steps are reduced (e.g., five steps in our main experiments), the attack cost of our evaluation method is comparable to the most widely used AutoAttack benchmark (Croce & Hein, 2020).
B.3 Additional Investigation on DiffPure with Our Evaluation
@@
Appendix C Derivation of Equations and Proofs of Theorems
C.1 The Derivation of Eq. 7
@@ Appendix C Derivation of Equations and Proofs of Theorems
Since ${\mathbf{x}}_{t}^{d}={\mathbf{x}}_{t}^{a}-k_{t}{\bm{\epsilon}}_{a}$ and ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, there is
${\mathbf{x}}_{t}^{d}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}.$ (13)
@@
$q({\mathbf{x}}_{t-1}^{d}|{\mathbf{x}}_{t}^{d},{\mathbf{x}}_{0})$ $=\frac{q({\mathbf{x}}_{t}^{d}|{\mathbf{x}}_{t-1}^{d},{\mathbf{x}}_{0})\cdot q({\mathbf{x}}_{t-1}^{d}|{\mathbf{x}}_{0})}{q({\mathbf{x}}_{t}^{d}|{\mathbf{x}}_{0})}\propto\exp(-\frac{1}{2}(\frac{({\mathbf{x}}_{t}^{d}-\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1}^{d}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\bm{\epsilon}}_{a})^{2}}{1-\alpha_{t}}$ (14)
$+\frac{({\mathbf{x}}_{t-1}^{d}-\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\bm{\epsilon}}_{a})^{2}}{1-\bar{\alpha}_{t-1}}-\frac{({\mathbf{x}}_{t}^{d}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a})^{2}}{1-\bar{\alpha}_{t}}))$
$=\exp\left(-\frac{1}{2}(A({\mathbf{x}}_{t-1}^{d})^{2}+B{\mathbf{x}}_{t-1}^{d}+C({\mathbf{x}}_{0},{\bm{\epsilon}}_{a},{\mathbf{x}}_{t}^{d}))\right),$
@@
$B=-2\sqrt{\alpha_{t}}\cdot\frac{{\mathbf{x}}_{t}^{d}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\bm{\epsilon}}_{a}}{1-\alpha_{t}}-2\frac{\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}+(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\bm{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}.$
In inference, we expect that ${\bm{\epsilon}}_{a}$ in Eq. 14 can be eliminated since only ${\mathbf{x}}_{0}^{a}$ is given by attackers and ${\bm{\epsilon}}_{a}$ cannot be decoupled from ${\mathbf{x}}_{0}^{a}$ directly. Based on the property of Gaussian distribution, eliminating all terms related to ${\bm{\epsilon}}_{a}$ in $B$ and $C$ in Eq. 5 can be achieved by eliminating all terms related to ${\bm{\epsilon}}_{a}$ in $B$. This yields:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\bm{\epsilon}}_{a}}{1-\alpha_{t}}-\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\bm{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}=0.$ (16)